OpenStudy (anonymous):

Billy has 1 gallon of paint. He is going to pour it into a paint tray that measures 10 inches wide, 14 inches long, and 4 cm deep. (1 gallon = 231 in3, 1 inch = 2.54 cm) Which of the following scenarios will happen? A. The paint will not fill the tray by 441 cm3. B. The paint will not fill the tray by 10.53 in3. C. The paint will fill the tray exactly. D. The paint will overfill the tray by 10.53 in3

OpenStudy (zepp):

I have detailed how to proceed to solve this question, why did you close that one and reposted it here? We don't seek answers here, we seek the path to the answers and understanding of these problems.

OpenStudy (zepp):

http://openstudy.com/study#/updates/53c41687e4b05c273e9cf87b Please re read it again, you can do it

OpenStudy (anonymous):

i was so lost

OpenStudy (zepp):

What you need to do here is to obtain the volume of the tray, and see if the paint could fill it Let's take for example the tray's volume is about 5 gallons and you have 10 gallons of paint, well, we can see that it will over fill by 5 gallons by doing PaintVolume - TrayVolume = 10 - 5 = 5 gallons overfill if we have on the other hand 3 gallons of paint, well, the tray will not fill completely 3-5 = -2, by 2 gallons

OpenStudy (triciaal):

first change the cm to inches to work with one unit of measure and it is better to work with inches since we already have the volume in inches. given 1 inch = 2.54 cm how many inches are the equivalent of 4 cm? the tray is like a box the volume = length * width* depth you have the measurements plug the numbers and find the volume of the tray now, Billy has 1 gallon of paint. you are given that 1 gallon = 231 in3 so you know the volume of one gallon of paint and you calculated the volume the tray can hold compare these amounts using the scenarios and decide which one is true. when considering option A remember to use the same unit. cm with cm or in with in